# Diffusion Model-Based Image Editing: A Survey

### v2

@@ 1 Introduction
In the realm of AI-generated content (AIGC), where artificial intelligence is leveraged to create and modify digital content, image editing is recognized as a significant area of innovation and practical applications. Different from image generation that creates new images from minimal inputs, image editing involves altering the appearance, structure, or content of an image, encompassing a range of changes from subtle adjustments to major transformations. This research is fundamental in various areas, including digital media, advertising, and scientific research, where altering visual content is essential. The evolution of image editing has reflected the advancements in digital technology, progressing from manual, labor-intensive processes to advanced digital techniques driven by learning-based algorithms. A pivotal advancement in this evolution is the introduction of Generative Adversarial Networks (GANs) [1, 2, 3, 4, 5, 6], which significantly enhance the possibilities for creative image manipulation. More recently, diffusion models have emerged in AIGC [7, 8, 9, 10, 11, 12, 13, 14, 1, 15], leading to notable breakthroughs in visual generation tasks. Diffusion models, inspired by principles from non-equilibrium thermodynamics [15], work by gradually adding noise to data and then learning to reverse this process from random noise until generating desired data matching the source data distribution. They can be roughly classified into denoising diffusion based [15, 16, 17, 18] and score-matching based [19, 20, 21, 22, 23]. Their adaptability and effectiveness have led to widespread applications across various tasks such as image generation [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38], video generation [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56], image restoration [57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71], and image editing.
@@
The application of diffusion models in image editing has experienced a surge in interest, evidenced by the significant number of research publications in this field over recent years. This growing attention highlights the potential and versatility of diffusion models in improving image editing performance compared to previous works. Given this significant advancement, it is essential to systematically review and summarize these contributions. However, existing survey literature on diffusion models concentrates on other specific visual tasks [72, 73, 74, 75], such as video applications [73] or image restoration and enhancement [74, 75]. Some surveys that mention image editing tend to provide only a cursory overview [76, 77, 78, 79, 80, 81, 82, 83], missing a detailed and focused exploration of the methods. To bridge this gap, we conduct a survey to offer an in-depth and comprehensive analysis that is distinctly focused on image editing. We delve deeply into the methodologies, input conditions, and a wide range of editing tasks achieved by diffusion models in this field. The survey critically reviews over 100 research papers, organizing them into three primary classes in terms of learning strategies: training-based approaches, testing-time finetuning approaches, and training and finetuning free approaches. Each class is further divided based on its core techniques, with detailed discussions presented in Sections 45, and 6, respectively. We also explore 10 distinct types of input conditions used in these methods, including text, mask, reference (Ref.) image, class, layout, pose, sketch, segmentation (Seg.) map, audio, and dragging points to show the adaptability of diffusion models in diverse image editing scenarios. Furthermore, our survey presents a new classification of image editing tasks into three broad categories, semantic editing, stylistic editing, and structural editing, covering 12 specific types. Fig. 1 visually represents the statistical distributions of research across learning strategies, input conditions, and editing task categories. In addition, we pay special attention to inpainting and outpainting, which together stand out as a unique type of editing. We explore both earlier traditional and current multimodal conditional methods, offering a comprehensive analysis of their methodologies in Section 7. We also introduce EditEval, a benchmark designed to evaluate text-guided image editing algorithms, as detailed in Section 8. In particular, an effective evaluation metric, LMM Score, is proposed by leveraging the advanced visual-language understanding capabilities of large multimodal models (LMMs). Finally, we present some current challenges and potential future trends as an outlook in Section 9.
@@ 2.1 Diffusion Models
Reverse Diffusion Process. The primary objective here is to learn the reverse of the forward diffusion process, with the aim of generating a distribution that closely aligns with the original unaltered data samples $\mathbf{z}_{0}$. In the context of image editing, $\mathbf{z}_{0}$ represents the edited image. Practically, this is achieved using a UNet architecture to learn a parameterized version of $p$. Given that the forward diffusion process is approximated by $q(\mathbf{z}_{T})\approx\mathcal{N}\left(\mathbf{0},\mathbf{I}\right)$, the formulation of the learnable transition is expressed as:
@@
DDIM Sampling and Inversion. When working with a real image $\mathbf{z}_{0}$, prevalent editing methods [84, 85] initially invert this $\mathbf{z}_{0}$ into a corresponding $\mathbf{z}_{T}$ utilizing a specific inversion scheme. Subsequently, the sampling begins from this $\mathbf{z}_{T}$, employing some editing strategy to produce an edited outcome $\tilde{\mathbf{z}}_{0}$. In an ideal scenario, direct sampling from $\mathbf{z}_{T}$, without any edits, should yield a $\tilde{\mathbf{z}}_{0}$ that closely resembles $\mathbf{z}_{0}$. A significant deviation of $\tilde{\mathbf{z}}_{0}$ from $\mathbf{z}_{0}$, termed reconstruction failure, indicates the inability of the edited image to maintain the integrity of unaltered regions in $\mathbf{z}_{0}$. Therefore, it is crucial to use an inversion method that ensures $\tilde{\mathbf{z}}_{0}\approx\mathbf{z}_{0}$.
The DDIM sampling equation [18] is:
@@ 2.2.1 Conditional Image Generation
While our primary focus is on diffusion models in image editing, it is important to acknowledge related areas like conditional image generation. Unlike image editing which involves altering certain parts of an existing image, conditional image generation is about creating new images from scratch, guided by specified conditions. Early efforts [32, 31, 87, 88, 89, 90] typically involve class-conditioned image generation, which usually incorporates the class-induced gradients via an additional pretrained classifier during sampling. However, Ho et al. [86] introduce the classifier-free guidance, which does not rely on an external classifier and allows for more versatile conditions, e.g., text, as guidance. Text-to-Image (T2I) Generation. GLIDE [34] is the first work that uses text to guide image generation directly from the high-dimensional pixel level, replacing the label in class-conditioned diffusion models. Similarly, Imagen [27] uses a cascaded framework to generate high-resolution images more efficiently in pixel space. A different line of research first projects the image into a lower-dimensional space and then applies diffusion models in this latent space. Representative works include Stable Diffusion (SD) [26], VQ-diffusion [91], and DALL-E 2 [25]. Following these pioneering studies, a large number of works [37, 92, 93, 94, 95, 96, 97] are proposed, advancing this field over the past two years. Additional Conditions. Beyond text, more specific conditions are also used to achieve higher fidelity and more precise control in image synthesis. GLIGEN [98] inserts a gated self-attention layer between the original self-attention and cross-attention layers in each block of a pretrained T2I diffusion model for generating images conditioned on grounding boxes. Make-A-Scene [99] and SpaText [100] use segmentation masks to guide the generation process. In addition to segmentation maps, ControlNet [101] can also incorporate other types of input, such as depth maps, normal maps, canny edges, pose, and sketches as conditions. Other methods like UniControlNet [102], UniControl [103], Composer [104], and T2I-Adapter [105] integrate diverse conditional inputs and include additional layers, enhancing the generative process controlled by these conditions. Customized Image Generation. Closely related to image editing within conditional image generation is the task of creating personalized images. This task focuses on generating images that maintain a certain identity, typically guided by a few reference images of the same subject. Two early approaches addressing this customized generation through few-shot images are Textual Inversion [106] and DreamBooth [107]. Specifically, Textual Inversion learns a unique identifier word to represent a new subject and incorporates this word into the text encoder’s dictionary. DreamBooth, on the other hand, binds a new, rare word with a specific subject by finetuning the entire Imagen [27] model with a few reference images. To combine multiple new concepts effectively, CustomDiffusion [108] only optimizes the cross-attention parameters in Stable Diffusion [26], representing new concepts and conducting joint training for the multiple-concept combination. Following these foundational works, subsequent methods [109, 110, 111, 112, 113, 114, 115, 116, 117] provide more refined control over the generated images, enhancing the precision and accuracy of the output.
@@ 2.2.2 Image Restoration and Enhancement
Image restoration (IR) stands as a pivotal task in low-level vision, aiming to enhance the quality of images contaminated by diverse degradations. Recent progress in diffusion models prompts researchers to investigate their potential for image restoration. Pioneering efforts integrate diffusion models into this task, surpassing previous GAN-based methods.
@@
Method
Condition(s)
Semantic Editing Stylistic Editing Structural Editing
Obj. Add.
Obj. Remo.
Obj. Repl.
Bg. Chg.
Emo. Expr. Mod.
Color Chg.
Style Chg.
Text. Chg.
Obj. Move.
Obj. Size. Chg.
Obj. Act. Chg.
Persp. /View. Chg.
DiffusionCLIP  [131]
Text, Class
Asyrp  [132]
Text, Class
EffDiff  [133]
Class
DiffStyler  [134]
Text, Class
StyleDiffusion  [135]
Ref. Image, Class
UNIT-DDPM  [136]
Class
CycleNet  [137]
Class
Diffusion Autoencoders  [138]
Class
HDAE  [139]
Class
EGSDE  [140]
Class
Pixel-Guided Diffusion  [141]
Seg. Map, Class
PbE  [142]
Ref. Image
RIC  [143]
Ref. Image, Sketch
ObjectStitch  [144]
Ref. Image
PhD  [145]
Ref. Image, Layout
DreamInpainter  [146]
Ref. Image, Mask, Text
Anydoor  [147]
Ref. Image, Mask
FADING  [148]
Text
PAIR Diffusion  [149]
Ref. Image, Text
SmartBrush  [150]
Text, Mask
IIR-Net  [151]
Text
PowerPaint  [152]
Text, Mask
Imagen Editor  [153]
Text, Mask
SmartMask  [154]
Text
Uni-paint  [155]
Text, Mask, Ref. Image
InstructPix2Pix  [156]
Text
MoEController  [157]
Text
FoI  [158]
Text
LOFIE  [159]
Text
InstructDiffusion  [160]
Text
Emu Edit  [161]
Text
DialogPaint  [162]
Text
Inst-Inpaint  [163]
Text
HIVE  [164]
Text
ImageBrush  [165]
Ref. Image, Seg. Map, Pose
InstructAny2Pix  [166]
Ref. Image, Audio, Text
MGIE  [167]
Text
SmartEdit  [168]
Text
iEdit  [169]
Text
TDIELR  [170]
Text
ChatFace  [171]
Text
UniTune  [172]
Text
Custom-Edit  [173]
Text, Ref. Image
KV-Inversion  [174]
Text
Null-Text Inversion  [175]
Text
DPL  [176]
Text
DiffusionDisentanglement  [177]
Text
Prompt Tuning Inversion  [178]
Text
@@
3 Categorization of Image Editing

### v3

@@ 1 Introduction
In the realm of AI-generated content (AIGC) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10], image generation and editing are recognized as significant for practical applications, including digital media, advertising, and scientific research. Distinct from image generation, which creates entirely new images from minimal inputs, image editing involves modifying the appearance, structure, or content of an existing image to achieve a desired outcome. This process can encompass a range of changes from subtle adjustments to moderate transformations, all without fundamentally altering the original image. The evolution of image editing has progressed from manual, labor-intensive processes to advanced learning-based algorithms. A pivotal advancement in this evolution is the introduction of Generative Adversarial Networks (GANs) [9, 11, 12, 13, 14, 15], which significantly enhance the possibilities for creative image manipulation.
More recently, diffusion models have emerged as a powerful tool in this domain. Inspired by principles from non-equilibrium thermodynamics [10], diffusion models work by gradually adding noise to data and then learning to reverse this process from random noise until generating desired data that matches the source data distribution [16, 17, 18, 19, 20, 21, 22, 23]. The application of diffusion models in image editing has garnered growing interest, as evidenced by the increasing number of research publications. This interest highlights the potential and versatility of diffusion models in improving image editing performance compared to previous methods. Given this significant advancement, it is essential to systematically review and summarize these contributions. However, despite this progress, there is a notable lack of surveys specifically dedicated to diffusion model-based image editing. Existing surveys [24, 25, 26, 27] on diffusion models tend to concentrate on other visual tasks, such as image generation [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43], restoration and enhancement [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58], or video generation [59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76]. Those that do mention image editing often provide only a cursory overview [77, 78, 79, 80, 81, 82, 83, 84], lacking a detailed and focused exploration of the methods. To address this gap, we conduct this survey to offer an in-depth and comprehensive analysis specifically focused on 2D image editing using diffusion models. Due to the extensive related literature, works that either focus on other relevant tasks or employ other techniques are not included as the primary review target.
@@
Specifically, we delve deeply into the methodologies, input conditions, and a wide range of editing tasks achieved by diffusion models in this field. The survey critically reviews over 100 research papers, organizing them into three primary classes—training-based approaches, testing-time finetuning approaches, and training and finetuning free approaches—based on whether their learning strategies require training, finetuning during inference, or can operate effectively without either. Each class is further divided based on its core techniques, with detailed discussions presented in Sections 45, and 6, respectively. We also explore 10 distinct types of input conditions used in these methods, including text, mask, reference (Ref.) image, class, layout, pose, sketch, segmentation (Seg.) map, audio, and dragging points to show the adaptability of diffusion models in diverse image editing scenarios. Furthermore, our survey presents a new classification of image editing tasks into three broad categories, semantic editing, stylistic editing, and structural editing, covering 12 specific types. Fig. 1 visually represents the statistical distributions of research across learning strategies, input conditions, and editing task categories. In addition, we pay special attention to inpainting and outpainting, which together stand out as a unique type of editing. We explore both earlier traditional and current multimodal conditional methods, offering a comprehensive analysis of their methodologies in Section 7. We also introduce EditEval, a benchmark designed to evaluate text-guided image editing algorithms, as detailed in Section 8. In particular, an effective evaluation metric, LMM Score, is proposed by leveraging the advanced visual-language understanding capabilities of large multimodal models (LMMs). Finally, we present some current challenges and potential future trends as an outlook in Section 9.
@@ 2.1 Diffusion Models
Reverse Diffusion Process. The reverse process, also called ancestral sampling [85], is defined by sampling a random Gaussian noise $\mathbf{z}_{T}$ and then for $t$ in $T,t-1,...,1$, sampling $\mathbf{z}_{t-1}$, until getting to $\mathbf{z}_{0}$. The primary objective here is to learn the reverse of the forward diffusion process, with the aim of generating a distribution that closely aligns with the original unaltered data samples $\mathbf{z}_{0}$. In the context of image editing, $\mathbf{z}_{0}$ represents the edited image. Practically, this is achieved using a UNet architecture to learn a parameterized version of $p$. Given that the forward diffusion process is approximated by $q(\mathbf{z}_{T})\approx\mathcal{N}\left(\mathbf{0},\mathbf{I}\right)$, the formulation of the learnable transition is expressed as:
@@
DDIM Sampling and Inversion. When working with a real image $\mathbf{z}_{0}$, prevalent editing methods [86, 87] initially invert this $\mathbf{z}_{0}$ into a corresponding $\mathbf{z}_{T}$ utilizing a specific inversion scheme. Subsequently, the sampling begins from this $\mathbf{z}_{T}$, employing some editing strategy to produce an edited outcome $\tilde{\mathbf{z}}_{0}$. In an ideal scenario, direct sampling from $\mathbf{z}_{T}$, without any edits, should yield a $\tilde{\mathbf{z}}_{0}$ that closely resembles $\mathbf{z}_{0}$. A significant deviation of $\tilde{\mathbf{z}}_{0}$ from $\mathbf{z}_{0}$, termed reconstruction failure, indicates the inability of the edited image to maintain the integrity of unaltered regions in $\mathbf{z}_{0}$. Therefore, using an inversion method that ensures $\tilde{\mathbf{z}}_{0}\approx\mathbf{z}_{0}$ is crucial. The DDIM sampling equation [18] is:
@@
2.2 Attention in Stable Diffusion
Cross-Attention in Stable Diffusion. In Stable Diffusion, cross-attention layers play a crucial role in fusing images and text, allowing T2I models to generate images that are consistent with textual descriptions. A cross-attention layer receives the query, key, and value matrices, i.e., $Q_{cross}$, $K_{cross}$, and $V_{cross}$, from the noisy image and prompt. Specifically, $Q_{cross}$ is derived from the spatial features of the noisy image by a linear layer, while $K_{cross}$ and $V_{cross}$ are projected from the textual embedding of the input prompt via a linear layer. The cross-attention map is defined as:
$M_{cross}=\text{Softmax}\left(\frac{Q_{cross}{K_{cross}}^{\mathsf{T}}}{\sqrt{d_{cross}}}\right),$ (13)
where $d_{cross}$ is the dimensionality of the keys and queries. The final output is defined as the fused feature of the text and image, denoted as $M_{cross}V_{cross}$. Intuitively, each cell $M_{cross}^{ij}$ in the cross-attention map determines the weights attributed to the value of the $j$-th text token relative to the spatial feature $i$ of the image. The cross-attention map enables the diffusion model to align the tokens of the prompt in the image area.
Self-Attention in Stable Diffusion. Unlike cross-attention, a self-attention layer receives the key matrix $K_{self}$ and the query matrix $Q_{self}$ from the noisy image through linear layers. The self-attention map is defined as:
$M_{self}=\text{Softmax}\left(\frac{Q_{self}{K_{self}}^{\mathsf{T}}}{\sqrt{d_{self}}}\right),$ (14)
where $d_{self}$ is the dimensionality of $K_{self}$ and $Q_{self}$. $M_{self}^{ij}$ of $M_{self}$ determines the weights assigned to the relevance of the $i$-th and $j$-th spatial features in the image and can affect the spatial layout and shape details of the generated image. Consequently, the self-attention map can be utilized to preserve the spatial structure characteristics of the original image during the editing process.
@@ 2.3 Related Tasks
While this survey focuses exclusively on diffusion model-based image editing, it is important to acknowledge related tasks that share similarities with image editing but differ in their objectives and methodologies. These tasks, such as conditional image generation, image restoration, and image composition, involve the creation, improvement, or compositing of images rather than the direct modification of existing content. Although these tasks are not the primary focus of this survey, understanding their distinctions and connections to image editing is crucial for a comprehensive view of how diffusion models are applied across various vision tasks.
@@ 2.3.1 Conditional Image Generation
Class-Conditioned Image Generation. Early efforts [36, 35, 88, 89, 90, 91] usually incorporate the class-induced gradients via an additional pretrained classifier during sampling. However, Ho et al. [32] introduce the classifier-free guidance, which does not rely on an external classifier and allows for more versatile conditions, e.g., text, as guidance.
Text-to-Image (T2I) Generation. GLIDE [38] is the first work that uses text to guide image generation directly from the high-dimensional pixel level, replacing the label in class-conditioned diffusion models. Similarly, Imagen [31] uses a cascaded framework to generate high-resolution images more efficiently in pixel space. A different line of research first projects the image into a lower-dimensional space and then applies diffusion models in this latent space. Representative works include Stable Diffusion (SD) [30], VQ-diffusion [92], and DALL-E 2 [29]. Following these pioneering studies, a large number of works [41, 93, 94, 95, 96, 97, 98] are proposed, advancing this field over the past two years.
Additional Conditions. Beyond text, more specific conditions are also used to achieve higher fidelity and more precise control in image synthesis. GLIGEN [99] inserts a gated self-attention layer between the original self-attention and cross-attention layers in each block of a pretrained T2I diffusion model for generating images conditioned on grounding boxes. Make-A-Scene [100] and SpaText [101] use segmentation masks to guide the generation process. In addition to segmentation maps, ControlNet [102] can also incorporate other types of input, such as depth maps, normal maps, canny edges, pose, and sketches as conditions. Other methods [103, 104, 105, 106, 107] integrate diverse conditional inputs and include additional layers, enhancing the generative process controlled by these conditions.
Personalized Image Generation. Closely related to image editing within conditional image generation is the task of creating personalized images. This task focuses on generating images that maintain a certain identity, typically guided by a few reference images of the same subject. Two early approaches are Textual Inversion [108] and DreamBooth [109]. Specifically, Textual Inversion learns a unique identifier word to represent a new subject and incorporates this word into the text encoder’s dictionary. DreamBooth binds a new rare word with a specific subject by finetuning the entire Imagen [31] model with a few reference images. For efficient training, some methods[110, 111, 112] depend on low-rank adaptation (LoRA) that focuses on adapting only a small subset of the model’s parameters. For example, Mix-of-Show adopts embedding-decomposed LoRA (ED-LoRA) for single-concept learning and gradient fusion for multi-concept fusion. Following these foundational works, subsequent methods [113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124] provide more refined control over the generated images, enhancing the precision and accuracy of the output.
Vector Graphics Generation. Scalable Vector Graphics (SVGs) can be scaled to any size, and are compact for digital icons, graphics, and stickers. VectorFusion [125] shows that a text-conditioned diffusion model trained on pixel representations of images can be used to generate SVG-exportable vector graphics. DiffSketcher [126] creates vectorized sketches by optimizing a set of Bézier curves with the score distillation sampling loss. A recent work [127] proposes a dual-domain (vector-pixel) diffusion with cross-modality impulse signals from each other.
@@ 2.3.2 Image Restoration and Enhancement
2.3.3 Image Composition
Image-Based Virtual Try-On. The goal of image-based virtual try-on (VTON) is to generate images of a person wearing a chosen garment while preserving the person’s identity and ensuring visual coherence [141, 142]. Traditional warp-based VTON methods [143, 144, 145, 146, 147] generally involve a two-step process: first, the garment is adjusted to match the target person’s pose, and then it is seamlessly overlaid onto the person’s image. Recently, with the rise of diffusion models, several VTON approaches [148, 149, 150, 151, 152, 153] have emerged, leading to significant improvements in generating realistic and high-quality try-on images. For example, TryOnDiffusion [149] unifies two UNets to preserve garment details and warp the garment for significant pose and body change in a single network. LaDIVTON [150] maps garment features to the CLIP embeddings to condition the latent diffusion model along with the warped input. StableVITON [151] employs a specialized zero cross-attention block to establish a semantic connection between the garment and the person’s features.
@@ 2.3.3 Image Composition
Method Condition(s) Semantic Editing Stylistic Editing Structural Editing
Obj. Add. Obj. Remo. Obj. Repl. Bg. Chg. Emo. Expr. Mod. Color Chg. Style Chg. Text. Chg. Obj. Move. Obj. Size. Chg. Obj. Act. Chg. Persp. /View. Chg.
DiffusionCLIP  [154] Text, Class
Asyrp  [155] Text, Class
EffDiff  [156] Class
DiffStyler  [157] Text, Class
StyleDiffusion  [158] Ref. Image, Class
UNIT-DDPM  [159] Class
CycleNet  [160] Class
Diffusion Autoencoders  [161] Class
HDAE  [162] Class
EGSDE  [163] Class
Pixel-Guided Diffusion  [164] Seg. Map, Class
PbE  [165] Ref. Image
RIC  [166] Ref. Image, Sketch
ObjectStitch  [167] Ref. Image
PhD  [168] Ref. Image, Layout
DreamInpainter  [169] Ref. Image, Mask, Text
Anydoor  [170] Ref. Image, Mask
FADING  [171] Text
PAIR Diffusion  [172] Ref. Image, Text
SmartBrush  [173] Text, Mask
IIR-Net  [174] Text
PowerPaint  [175] Text, Mask
Imagen Editor  [176] Text, Mask
SmartMask  [177] Text
Uni-paint  [178] Text, Mask, Ref. Image
InstructPix2Pix  [179] Text
MoEController  [180] Text
FoI  [181] Text
LOFIE  [182] Text
InstructDiffusion  [183] Text
Emu Edit  [184] Text
DialogPaint  [185] Text
Inst-Inpaint  [186] Text
HIVE  [187] Text
ImageBrush  [188] Ref. Image, Seg. Map, Pose
InstructAny2Pix  [189] Ref. Image, Audio, Text
MGIE  [190] Text
SmartEdit  [191] Text
iEdit  [192] Text
TDIELR  [193] Text
ChatFace  [194] Text
UniTune  [195] Text
Custom-Edit  [196] Text, Ref. Image
KV-Inversion  [197] Text
Null-Text Inversion  [198] Text
DPL  [199] Text
DiffusionDisentanglement  [200] Text
Prompt Tuning Inversion  [201] Text
@@
Image-Guided Object Composition. Image-guided object composition [167, 165, 238, 259, 260] incorporates specific objects and scenarios from user-provided photos, potentially with the assistance of a text prompt. Recent studies [167, 165] effectively incorporate additional guiding images into diffusion models through retraining pre-trained models on their datasets. TF-ICON [238] involves the gradual injection of composite self-attention maps through multiple samplers. PrimeComposer [261] formulates this task as a subject-guided local editing problem and depicts the object through attention steering.
3 Scope and Categorization